A dart board has a diameter of 40 cm and is divided into 20 congruent
sectors. What is the area of one of the sectors?. please help me answer this

Answer by Theo(13342)

the radius is equal to half the diameter = 20 cm.
the area of the circle formed is equal to pi * 20^2 = 400 * pi
there are 20 sectors that are all congruent to each other, so the area of each individual sector is 400 * pi / 20 = 20 * pi.

you can also calculate it another way.
if there are 20 congruent sectors, then each sector has a central angle of 360 / 20 = 18.
the area of the sector is equal to the central angle of the sector divided by 360 * the area of the circle.
18/360 * 400 * pi = 1/20 * 400 * pi = 400 / 20 * pi = 20 * pi.